### v1

# GSR-Bench: A Benchmark for Grounded Spatial Reasoning Evaluation via Multimodal LLMs

Abstract
The ability to understand and reason about spatial relationships between objects in images is an important component of visual reasoning. This skill rests on the ability to recognize and localize objects of interest and determine their spatial relation. Early vision and language models (VLMs) have been shown to struggle to recognize spatial relations. We extend the previously released What’sUp dataset Kamath et al. (2023) and propose a novel comprehensive evaluation for spatial relationship understanding that highlights the strengths and weaknesses of 27 different models. In addition to the VLMs evaluated in What’sUp, our extensive evaluation encompasses 3 classes of Multimodal LLMs (MLLMs) that vary in their parameter sizes (ranging from 7B to 110B), training/instruction-tuning methods, and visual resolution to benchmark their performances and scrutinize the scaling laws in this task.
GSR-Bench: A Benchmark for Grounded Spatial Reasoning Evaluation
via Multimodal LLMs
Navid Rajabi George Mason University nrajabi@gmu.edu                        Jana Košecká George Mason University kosecka@gmu.edu
1 Introduction
Earlier efforts for benchmarking vision and language models (VLMs) were developed for cross-modal and/or dual-encoder, end-to-end models, like LXMERT Tan and Bansal (2019), CLIP Radford et al. (2021), BLIP Li et al. (2022), with the focus on downstream tasks performances such as VQA Antol et al. (2015), GQA Hudson and Manning (2019), referring expressions Kazemzadeh et al. (2014), image-text matching or image/text retrieval.
Figure 1: LLaMA-3-LLaVA-NeXT-8B achieves the overall accuracy of 86.1%, compared to 60.4% by XVLM-COCO, in What’sUp benchmark, reaching the best trade-off between accuracy and parameters size, since it performs only 1.1% lower than LLaVA-NeXT-34B, which has $\times$4.25 number of parameters.
Figure 2: Our pipeline overview for spatial relationship understanding prompting, shown in the top two figures, and our depth-augmented prompting, shown in the bottom figure.
While spatial relations are often part of VQA datasets, the evaluation of spatial reasoning is often conflated with grounding referring expressions or objects and their attributes [^1] VQA example question may be: “Is there a woman to the left of the person that is wearing a wetsuit?”. To isolate these issues, authors in Kamath et al. (2023) introduced a new benchmark that focuses on spatial relationship understanding only. Using image-text matching evaluation methodology, they showed that contrastive models such as CLIP, BLIP, and their follow-up variants struggle to understand spatial relations with the best accuracy around $61$%.
Recent advances in generative large multi-modal models have shown remarkable visual knowledge and reasoning capabilities. We revisit the spatial relationship understanding in the context of MLLMs and extend the existing What’sUp benchmark Kamath et al. (2023) to include bounding box annotations and depth information. Compositional spatial relationship understanding requires successful recognition of objects and determining their locations. Furthermore, the knowledge of scene depth helps to disambiguate certain relationships (e.g., "in front of" or "behind"). The availability of this information can support a grounded understanding of spatial relations and will contribute to the fine-grained evaluation of large generative MLLMs, which lag behind their earlier counterparts. A few exceptions are multi-task multi-modal benchmarks like MMBench Liu et al. (2023d) and its related benchmarks that focus on evaluating several MLLMs for both visual recognition tasks and description generation. Given the simple structure of spatial clauses, we can study separately the ability of the model to ground the subject and object in the clause, and the effect and means of incorporating the depth information. The contributions of this work can be summarized as follows:
  • Extended What’sUp spatial relationship dataset with depth, masks, and bounding box annotations.
  • Design of different prompting strategies through structured prompting for the evaluation of grounding and spatial reasoning.
  • Comprehensive evaluation and comparison of 18 VLMs and 9 MLLMs, with various sizes, resolutions, pre-training/instruction-tuning, and prompting strategies.
Model Num Subset A Subset B COCO-spatial GQA-spatial Total Average
Params Sub-Obj Sub-Obj One-Obj Two-Obj One-Obj Two-Obj
CLIP ViT-B/32 Radford et al. (2021) 151M 30.3 31.6 43.7 51.1 46.5 47.4 41.8
CLIP ViT-L/14 428M 26.5 25.7 49.2 49.8 46.1 48.5 41.0
NegCLIP Yuksekgonul et al. (2022) 32.5 36.3 47.4 46.4 45.3 46.7 42.4
RoBERTaCLIP Kamath et al. (2023) 25.2 25.0 46.3 53.6 50.8 48.8 41.6
CoCa Yu et al. (2022) 2.1B 29.4 29.4 48.1 45.2 45.0 49.1 41.0
XVLM 4M Zeng et al. (2021) 216M 40.0 23.0 58.4 65.0 62.8 54.6 50.6
XVLM 16M 216M 50.7 33.1 65.4 64.5 63.2 53.3 55.0
BLIP 14M Li et al. (2022) 583M 38.8 38.2 54.2 53.9 49.1 50.5 47.5
BLIP 129M 583M 30.3 30.4 44.8 53.9 50.5 47.4 42.9
BLIP2-ITM Li et al. (2023) 188M 44.9 30.4 48.3 57.7 46.0 53.6 46.8
BLIP2-ITC 188M 35.9 22.1 55.6 51.8 52.6 49.5 44.6
FLAVA Singh et al. (2022) 33.7 27.2 50.3 55.0 52.2 51.2 44.9
CoCa-Caption 2.1B 25.5 22.8 45.9 51.4 48.5 50.5 40.8
XVLM-Flickr30K 216M 45.1 43.4 63.1 67.3 64.7 58.1 56.9
XVLM-COCO 216M 41.7 42.4 68.4 73.6 69.1 67.0 60.4
BLIP-Flickr30K 583M 29.6 38.0 50.0 58.4 50.3 47.4 45.6
BLIP-COCO 583M 35.7 29.9 46.4 56.4 50.3 52.6 45.2
BLIP-VQA 583M 57.8 37.7 63.6 60.5 63.8 52.9 56.0
LLaVA-1.5-Vicuna 7B 25.0 31.9 90.4 66.6 91.2 62.9 61.3
LLaVA-1.5-Vicuna 13B 58.5 28.2 92.5 78.9 93.1 82.8 72.3
LLaVA-NeXT-Mistral 7B 37.4 22.0 81.1 60.4 89.4 57.0 57.9
LLaVA-NeXT-Vicuna 7B 38.6 26.2 95.5 71.8 97.6 79.0 68.1
LLaVA-NeXT-Vicuna 13B 75.0 20.1 95.6 78.6 97.6 84.9 75.3
LLaMA-3-LLaVA-NeXT 8B 94.2 60.8 95.1 83.9 97.8 85.2 86.1
LLaVA-NeXT-Yi 34B 82.3 75.7 94.8 87.7 91.5 91.1 87.2
LLaVA-NeXT-Qwen1.5 110B 93.9 54.2 90.6 84.1 96.2 94.2 85.4
Intern-VL-Chat-1.5 26B 92.2 61.8 95.1 82.3 97.8 82.8 85.3
Random Chance 25.0 25.0 50.0 50.0 50.0 50.0 41.7
Table 1: Template-based generation (TG) results using CircularEval. The first two sections come from What’sUp Kamath et al. (2023) results. The rest shows our LLaVA 1.5, 1.6, and InternVL-1.5 prompting results. Our best-performing is shown in bold, 2nd-best with underline, and What’sUp best-performing with italic underline.
2 GSR Benchmark
We extend carefully curated What’sUp dataset Kamath et al. (2023) that is comprised of Subset A containing pairs of objects in unambiguous spatial relations, being "on", "under", "left of" or "right of" a table, chair, or armchair, and Subset B containing an object "in front of", "behind", "left" or "right" of another object on a tabletop, and subsets of COCO-Spatial and GQA-Spatial with either one or two objects occurring, accompanied by spatial clauses like "on top of", "on the bottom of", "right of", or "left of". To study the grounding in this context, we annotate the dataset with bounding box coordinates and segmentation masks for all the objects mentioned in the captions and the depth maps for the images. We leverage GroundingDINO Liu et al. (2023c) as an open-vocabulary object detector, Segment Anything (SAM) Kirillov et al. (2023) for the object mask segmentation, and ZoeDepth Bhat et al. (2023) for monocular depth estimation. In the next section, we explain in detail how these additional annotations enable a more rigorous and grounded evaluation of spatial reasoning and its components [^2] All the code and data will be publicly available..
Model Subset A Subset B COCO-spatial GQA-spatial Avg G-Score
Sub Obj Sub Obj One-Obj Sub Obj One-Obj Sub Obj
LLaVA-1.5-Vicuna-7B 9.7 79.4 51.5 25.7 47.4 49.8 48.0 31.9 55.0 47.8 44.62
LLaVA-1.5-Vicuna-13B 13.8 86.1 77.4 32.3 60.9 61.8 61.0 42.1 72.2 59.8 56.74
LLaVA-NeXT-Vicuna-7B 14.1 99.0 95.8 66.7 81.9 84.5 77.7 45.5 60.1 56.0 68.13
LLaVA-NeXT-Mistral-7B 13.1 82.3 93.9 60.0 87.1 86.8 85.7 69.2 85.9 81.8 74.58
LLaVA-NeXT-Vicuna-13B 15.3 84.0 95.3 67.6 87.1 90.2 83.9 69.6 85.9 80.4 75.93
LLaMA-3-LLaVA-NeXT-8B 19.2 99.3 96.6 73.8 85.7 87.5 83.2 69.0 84.5 80.4 77.92
LLaVA-NeXT-Yi-34B 21.1 100.0 97.8 78.9 83.7 85.7 81.4 70.0 88.0 83.5 79.01
LLaVA-NeXT-Qwen1.5-110B 29.4 98.5 98.8 80.1 88.7 88.2 86.4 74.9 86.9 84.2 81.61
GroundingDINO [avg($\rho$)] 58.8 92.0 78.1 70.1 62.3 62.8 59.3 59.4 70.4 65.2 67.84
GroundingDINO [$\Sigma$($\rho$geq 0.5)/$t$] 68.9 100.0 90.0 88.7 71.0 73.6 66.4 59.1 76.3 71.1 76.51
Table 2: Grounding/Localization results. AVG G-Score refers to the mean accuracy of IoU $\geq$ 0.5. The bottom two rows refer to the GroundingDINO mean confidence scores ($\rho$), and mean accuracy of $\rho$geq 0.5, respectively.
3 GSR-Bench Experiments
Grounded spatial reasoning evaluation is typically done using image-text matching, binary VQA, or multiple-choice VQA. Further evaluations include subject and/or object grounding and localization; and exploring the effect of using depth information. In addition to 18 VLMs that have been evaluated in Kamath et al. (2023), we focus on the probing of open-source generative MLLMs like LLaVA and InternVL [^3] InternVL is the leading model in MMBench. using structured generation methodologies of Multiple choice (MC) and Template-based generation (TG). In MC prompting, captions for each image are represented as A, B, C, and D options for Subset A and Subset B, while A and B options for COCO-Spatial and GQA-Spatial Subsets. Then, the model is prompted to choose the correct letter as the final answer. In TG prompting, as shown in Figure 3, we append the correct format of the entire caption to the prompt, in which the spatial clause acts as the placeholder for the correct spatial relation option. In this way, we are able to leverage LLMs’ open-ended generation capability, handle the models’ verbosity by enforcing the correct answer structure, and overcome the biases observed in MC prompting simultaneously (See Figure 2).
Figure 3: TG sample prompt structure.
We ran each prompt with 4 different permutations so as to vary the position of the answer among the choices in MC and the list of options in TG prompting. An instance is considered correct if all four options are predicted correctly, known as CircularEval, introduced in MMBench Liu et al. (2023d). As opposed to the CircularEval, there exists VanillaEval, which only asks the model to choose the correct answer from a list of options once and has been shown to be prone to bias in recent studies. We first ran our experiments using MC prompting and observed a significant degree of bias among the models when the position of the answer varied among the choices of A, B, C, or D. This bias and sensitivity turned out to be even more detrimental in smaller models, while larger models like LLaVA-NeXT-Yi-34B and LLaMA-3-LLaVA-NeXT-8B showed significantly higher robustness (See Figure 4 in the Appendix for details). This phenomenon also corroborates the findings of multiple recent studies in LLMs Zheng et al. (2023); Pezeshkpour and Hruschka (2023); Wang et al. (2023); Xue et al. (2024); Wang et al. (2024). According to this observation, we opted for TG prompting, accompanied by the CircularEval methodology, inspired by Gemini 1.5 Pro Reid et al. (2024). See Table 1 for the TG prompting results, where rows in section 1 and 2 come from the What’sUp benchmark Kamath et al. (2023), section 3 refers to LLaVA-1.5 models Liu et al. (2023b), section 4 to the LLaVA-NeXT models Li et al. (2024); Liu et al. (2024), and section 5 to the InternVL-1.5 results Chen et al. (2024).
Grounding/Localization Evaluation.
[…]
Model
w/o
depth
with
depth
InternVL-Chat-1.5-26B 26.5 40.7
LLaMA-3-LLaVA-NeXT-8B 53.4 60.3
LLaVA-NeXT-Yi-34B 64.7 81.9
Table 3: DAP results for behind & in front of cases.
Depth-Augmented Prompting (DAP).
The experiments in Table 1 revealed that Subset B is the lowest-performing, with many instances requiring reasoning about "behind" and "in front of" spatial clauses. We propose to incorporate the depth values of subject and object into the prompt, as a hint to the model, utilizing our augmented benchmark annotations, depicted in Figure 2. We show that this minimal change improves the accuracy of top-3 performing models in these instances of Subset B, reported by CircularEval in Table 3.
4 Discussion
According to Table 1 and 2, there is a positive correlation, even stronger in grounding, between scaling the LLM size & visual resolution, and the overall accuracy in both tasks. Conversely, there exist multiple exceptions, which are inevitable to concretely justify due to various intervening factors, such as (1) differences in training/fine-tuning & architectures and (2) release date and further instruction-tuning of the LLMs, like LLaMA-3-8B, which has the most-recent knowledge cut-off.
Grounding small objects, which refers to the Sub column in Subset A, seems challenging for all, and worst in smaller models, according to Table 2. We also observed a plateau in Table 1, especially in Qwen-1.5-110B, which is the largest ever released open-source MLLM at the moment. This could be a sign of saturation where the reasoning capability flattens out, although scaling still improves grounding, shown in Table 2.
5 Conclusions
In this work, we introduce a new benchmark for grounded spatial reasoning by enriching the What’sUp dataset with additional supervision for a more fine-grained assessment of MLLM’s spatial understanding. We also propose a new compositional evaluation methodology for (1) a stricter assessment of spatial relationship understanding through CircularEval, and (2) measuring the model’s grounding capability using the labels we generate through our cost-effective auto-annotation pipeline. Our evaluations reveal the superiority of LLaVA MLLMs over the best-performing VLMs evaluated in What’sUp, like XVLM, by a significant margin of $\sim$+∼ +26.8%. Future works may investigate the remaining gap between the top open-source MLLMs and human-level accuracy.
Limitations
Small-scale Dataset:
Our split sizes remain the same as the What’sUp dataset in which Subset A has 412, Subset B has 408, COCO-Spatial-One has 2247, COCO-Spatial-Two has 440, GQA-Spatial-One has 1160, and GQA-Spatial-Two has 291 instances. Although this benchmark includes 4,958 image instances in total, each instance covering one or two objects, with various domain shifts in each 6 split, it is smaller than already existing benchmarks related to spatial reasoning, like Visual Genome Krishna et al. (2017), GQA Hudson and Manning (2019), VSR Liu et al. (2022), SpatialSense Yang et al. (2019), MMBench Liu et al. (2023d), etc. The reason is that this work aims to provide a carefully curated benchmark for spatial relationship understanding evaluation in a controlled setting to abstract away intervening factors that make the evaluations noisy.
Limited Spatial Prepositions:
Following the What’sUp dataset, our benchmark is also confined to the primitive spatial clauses of on, under, behind, in front of, to the left of, to the right of, below and above, when having two objects involved in the caption, and, on the top, on the bottom, on the left and on the right when having only one object in the caption, like in COCO-Spatial-One and GQA-Spatial-One.
Lack of Robustness in MC Prompting:
In addition to the similar findings of MC noisiness in LLMs that we discussed earlier, we hypothesize that the higher degree of variance in multiple-choice results in the last two subsets (COCO and GQA), which is more significant in the smaller models, could be due to the language domain distribution shift. Most of the LLMs and MLLMs are being trained and evaluated with 4 options in the multiple-choice settings. Conversely, in the last two subsets, we have two captions per image, which means we only provide options A and B to the model in the prompt instead of ABCD without any fine-tuning for this task or this specific type of prompting.
Intern-VL-1.5 Poor Grounding Observation:
An unexpected, significant noisiness in the output of grounding/localization prompting of InternVL-1.5 model prevented us from analyzing and reporting the results for this model, which requires further investigation since a similar behavior has been observed through our interaction with the InternVL-1.5 demo, as well.
Depth Augmentation Nuances:
The issue we noticed in the DAP experiment was the distraction the depth hint can cause in cases where multiple correct relationships hold in the image. For instance, object A can be to the left of object B, and also in front of object B, at the same time. So, in these ambiguous cases, incorporating depth could make the model’s decision biased towards the in front of preposition, while the ground-truth might be to the left of in this case. Therefore, we believe that trying both prompts, with and w/o depth hint, would be helpful for disambiguation in such cases.
No Human Annotation:
Due to the resource constraints, our extended benchmark relies on the pseudo-labels we generate using state-of-the-art, off-the-shelf models like GroundingDINO, SAM, and ZoeDepth. Future works could incorporate human inspection and labeling for further robustness in annotations.
Appendix A Appendix
The appendix is organized as follows:
  • Figure 4 demonstrates the biases of multiple-choice (MC) prompting.
  • Figures 5 - 11 depict the distributions of objects occurring in the captions.
  • Figure 12 shows sample failures in grounding small objects in Subset A.
Figure 4: Sensitivity of the models to different permutations of choice order, in the multiple-choice (MC) experiment, which is more significant in the smaller models, and when having two choices of A and B instead of regular 4-choice of A, B, C, and D. LLaVA-NeXT-Yi-34B demonstrates an excellent robustness against this issue.
Figure 5: Subset A - subjects and objects distributions.
Figure 6: Subset B - subjects and objects distributions.
Figure 7: COCO Spatial One - objects distribution.
Figure 8: COCO Spatial Two - subjects distribution.
Figure 9: COCO Spatial Two - objects distribution.
Figure 10: GQA Spatial Two - subjects distribution.
Figure 11: GQA Spatial Two - objects distribution.
(a) toy cactus on chair
(b) wineglass under armchair
(c) tape under armchair
(d) sunglasses on table
(e) spatula on chair
(f) remote on armchair
(g) orange right of armchair
(h) ball of yarn left of table
(i) wineglass on table
(j) banjo under armchair
Figure 12: Sample failures in small objects grounding (i.e., IoU $<0.5$), which refers to the Sub column results of Subset A in Table 2. The pseudo-ground-truth bounding box, which is the GroudningDINO output, is indicated in green, and the output of LLaVA-NeXT-Qwen-1.5-110B, which is the best-performing MLLM in our grounding/localization experiment, is demonstrated in yellow.